# arXiv:2603.01962v1[quant-ph] 02 Mar 2026

### v1

@@ Abstract
On the other hand, the dynamic Bayesian network method is based on projective measurements in the eigenbasis of $\rho_{\textrm{i},\textrm{f}}$, instead of that of $H_{\textrm{i},\textrm{f}}$, hence preserving initial and final coherences in the energy basis lutz_2020 ; par20 ; mic20a ; str20 ; zha22 ; lutz_2023 ; lutz_2024 . The joint probability $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})$ of the unmeasured energy eigenvalues is then obtained using Bayesian inference. Let us concretely write the eigendecomposition of the initial and final states as $\rho_{\textrm{i},\textrm{f}}=\sum_{s}\lambda_{\textrm{i},\textrm{f}}^{s}P_{\textrm{i},\textrm{f}}^{s}$ where $P_{\textrm{i},\textrm{f}}^{s}$ is the eigenprojector corresponding to the eigenvalue $\lambda_{\textrm{i},\textrm{f}}^{s}$. By now performing the first projective measurement on one half of the measurement ensemble and the second projective measurement on the remaining, time evolved, half of the ensemble, the joint distribution of measurement outcomes is $p(\lambda_{\mathrm{i}}^{s},\lambda_{\mathrm{f}}^{q})=\text{tr}[P_{\textrm{f}}^{q}\,\mathrm{M}(P_{\textrm{i}}^{s}\rho_{\textrm{i}}P_{\textrm{i}}^{s})]$. The next step is to evaluate the posterior conditional probabilities for each Hamiltonian eigenvalue using the standard Born rule, $p(e_{\mathrm{i}}^{j}|\lambda_{\mathrm{i}}^{s})=\text{tr}[\Pi_{\textrm{i}}^{j}P_{\textrm{i}}^{s}\rho_{\textrm{i}}P_{\textrm{i}}^{s}/\text{tr}(P_{\textrm{i}}^{s}\rho_{\textrm{i}})]$ and $p(e_{\mathrm{f}}^{k}|\lambda_{\mathrm{f}}^{q})=\text{tr}[\Pi_{\textrm{f}}^{k}P_{\textrm{f}}^{q}\rho_{\textrm{f}}P_{\textrm{f}}^{q}/\text{tr}(P_{\textrm{f}}^{q}\rho_{\textrm{f}})]$. Summing over all possible measurement outcomes, one finds lutz_2020 ; par20 ; mic20a ; str20 ; zha22 ; lutz_2023 ; lutz_2024 , $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})=\sum_{s,q}p(e_{\mathrm{i}}^{j}|\lambda_{\mathrm{i}}^{s})\,p(e_{\mathrm{f}}^{k}|\lambda_{\mathrm{f}}^{q})\,p(\lambda_{\mathrm{i}}^{s},\lambda_{\mathrm{f}}^{q})$. The two joint distributions, $p_{\textrm{TPM}}$ and $p_{\textrm{DBN}}$, coincide in the absence of quantum coherence in the energy basis, yielding identical energy statistics. However, because the states $\rho_{\textrm{i},\textrm{f}}$ are diagonal in their eigenbases, measurement backaction is strongly reduced in the DBN protocol in the presence of energy coherences, compared to the TPM approach. Importantly, the Bayesian network distribution $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})$ is accessible experimentally mic20a .
@@
The joint probability $p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})$ in Eq. (S1) can be derived for the quantum Otto engine in a similar manner by extending the previous considerations to the four corners of the cycle. The explicit (and lengthy) expressions for the work output distribution $P(w)$ for the two measurement schemes are given in the Supplemental Material. We here indicate the corresponding average work outputs $\langle w\rangle=\int dw\,w\,P(w)$. We have
@@
Remarkably, the average work, $\langle w\rangle_{\mathrm{DBN}}$ in Eq. (3), obtained via Bayesian networks always agrees exactly with the unmeasured result $W$. By contrast, the two-point measurement expression, $\langle w\rangle_{\mathrm{TPM}}$ in Eq. (2), differs from it in general due to measurement backaction. A comparison of Eqs. (2) and (3) reveals that this discrepancy originates from the presence of quantum coherence during the cycle: $\langle w\rangle_{\mathrm{TPM}}=\langle w\rangle_{\mathrm{DBN}}=W$ only if the states of the working substance after the hot and cold isochores are diagonal in the eigenbasis of their respective Hamiltonians, that is, $\tilde{\rho}_{1}=\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})$ after the cold isochore and $\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})=\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})$ after the hot isochore. For the specific Hamiltonian considered here, this is achieved either when coherences are not dynamically induced ($g_{i}=0$), or when the system fully thermalizes with the baths ($t_{\mathrm{h},\mathrm{c}}\to\infty$). In these cases the two work distributions (1) also coincide, $P_{\mathrm{TPM}}(w)=P_{\mathrm{DBN}}(w)$. For arbitrary system Hamiltonians $H_{i}$, and finite thermalization times, these conditions can be generically formulated as (Supplemental Material)
$\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})\!=\mathrm{D}_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})\text{ and }\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}(\tilde{\rho}_{1})\!=\mathrm{D}_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}(\tilde{\rho}_{1})\!\!$ (4)
since the output of the generalized amplitude damping map $\mathcal{T}_{t}$ is diagonal only if the input is already diagonal. It is worth noting that dynamic Bayesian networks are different from so-called collective measurements which have been shown to reduce measurement backaction acin_2017 ; wu19 . Whereas the measured average work in that scheme approaches the unmeasured value, after proper parameter optimization, it does not reach it in general.
@@
Both measurement protocols are unavoidably accompanied by measurement backaction. However, the above findings suggest that the backaction induced by the dynamic Bayesian network scheme is significantly weaker than that of the two-point measurement approach. To quantify the disturbance induced by a measurement sequence, we compare the unmeasured steady state $\tilde{\rho}_{1}$ with the averaged state obtained after a full engine cycle interspersed by five measurements, back to the initial corner of the cycle. We find $\langle\rho\rangle_{\textrm{TPM}}=\mathrm{D}_{\mathrm{e}}\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathrm{D}_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathrm{D}_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})\neq\tilde{\rho}_{1}$ whereas $\langle\rho\rangle_{\textrm{DBN}}=\tilde{\rho}_{1}$. The Bayesian network approach therefore does not perturb the state of the working substance on average. In this sense, it is minimally invasive, in sharp contrast to the two-point measurement protocol. This difference arises because, in the Bayesian network scheme, measurements are performed in the eigenbasis of the density operator, in which the state is trivially diagonal. Figure S3 displays the trace distance, $d(\langle\rho\rangle,\tilde{\rho}_{1})$ nie10 , between the averaged measured state and the unmeasured state, as a function of the thermalization rate $\lambda$. These results indicate that two-point measurement experiments become increasingly invasive as the amount of coherence generated during the engine cycle grows for smaller $\lambda$.
@@
Equation (5) was derived assuming full thermalization, that is, in the absence of quantum coherence agarwalla_2021_2 ; agarwalla_2021 ; das23 ; moh23 ; moh23a ; watanabe_2025 . Figure 4 reveals that the bound is actually violated in the presence of quantum coherence, and this for both measurement schemes, for finite relaxation times. Coherence is seen to augment work output fluctuations. We stress that the value of the transverse driving strength is here one order of magnitude smaller than the frequencies to ensure that the engine runs in the quasistatic regime.
@@ I.2 Dynamic Bayesian network approach (DBN)
In the case of the DBN protocol lutz_2020s ; par20s ; mic20as ; str20s ; zha22s ; lutz_2023s ; lutz_2024s , it is convenient to consider five copies of the working substance lutz_2023s : $\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}$ (If only $p_{w}$ and $p_{q_{\mathrm{h}}}$ are required, only four copies are necessary). Given that the quantum engine reaches a steady state after a few cycles, in order to obtain the five copies, it is enough to initialize the engine with five arbitrary states $\rho$ and wait until they all reach the steady state $\tilde{\rho}_{1}$. Let $\tilde{\rho}_{i}=\sum_{\alpha}\lambda_{i}^{\alpha}P_{i}^{\alpha}$ be the spectral decomposition of the state of the working substance at corner $i$ of the engine cycle. The measurement protocol to obtain $p^{\mathrm{DBN}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ proceeds as follows:
@@
    $P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}\otimes\mathcal{U}_{\mathrm{k}}[P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}/\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})]\otimes\tilde{\rho}_{2}\otimes\tilde{\rho}_{2}\otimes\tilde{\rho}_{2}.$ (S18)
@@
    $P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}\otimes P_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}/\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})]P_{2}^{\beta}\otimes\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}/\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})]\otimes\tilde{\rho}_{3}\otimes\tilde{\rho}_{3}.$ (S19)
@@
    $P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}\otimes P_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}/\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})]P_{2}^{\beta}\otimes P_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}/\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})]P_{3}^{\gamma}\otimes\mathcal{U}_{\mathrm{e}}[P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}/\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})]\otimes\tilde{\rho}_{4}.$ (S20)
@@
    $P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}\otimes P_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}/\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})]P_{2}^{\beta}\otimes P_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}/\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})]P_{3}^{\gamma}$ (S21)
    $\otimes P_{4}^{\delta}\mathcal{U}_{\mathrm{e}}[P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}/\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})]P_{4}^{\delta}\otimes\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}[P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}/\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})].$
@@
    $P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}\otimes P_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}/\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})]P_{2}^{\beta}\otimes P_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}/\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})]P_{3}^{\gamma}$ (S22)
    $\otimes P_{4}^{\delta}\mathcal{U}_{\mathrm{e}}[P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}/\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})]P_{4}^{\delta}\otimes P_{1}^{\epsilon}\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}[P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}/\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})]P_{1}^{\epsilon}.$
@@
$p(\lambda_{1}^{\alpha},\lambda_{2}^{\beta},\lambda_{3}^{\gamma},\lambda_{4}^{\delta},\lambda_{1}^{\epsilon})$ $=\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{P_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{P_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}$ (S23)
$\times\text{tr}\left\{P_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}\,\text{tr}\left\{P_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left[\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right]\right\}.$

### v2

# Minimal-backaction work statistics of coherent engines

@@ Abstract
On the other hand, the dynamic Bayesian network method is based on projective measurements in the eigenbasis of $\rho_{\textrm{i},\textrm{f}}$, instead of that of $H_{\textrm{i},\textrm{f}}$, hence preserving initial and final coherences in the energy basis lutz_2020 ; par20 ; mic20a ; str20 ; zha22 ; lutz_2023 ; lutz_2024 . The joint probability $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})$ of the unmeasured energy eigenvalues is then obtained using Bayesian inference. Let us concretely write the eigendecomposition of the initial and final states as $\rho_{\textrm{i},\textrm{f}}=\sum_{\alpha}\lambda_{\textrm{i},\textrm{f}}^{\alpha}\mathrm{P}_{\textrm{i},\textrm{f}}^{\alpha}$ where $\mathrm{P}_{\textrm{i},\textrm{f}}^{\alpha}$ is the eigenprojector corresponding to the eigenvalue $\lambda_{\textrm{i},\textrm{f}}^{\alpha}$. By now performing the first projective measurement on the complete measurement ensemble, dividing it into two halves, time evolving one of them and then projectively measuring it, the joint distribution of measurement outcomes is $p(\lambda_{\mathrm{i}}^{\alpha},\lambda_{\mathrm{f}}^{\beta})=\text{tr}[\mathrm{P}_{\textrm{f}}^{\beta}\,\mathrm{M}(\mathrm{P}_{\textrm{i}}^{\alpha}\rho_{\textrm{i}}\mathrm{P}_{\textrm{i}}^{\alpha})]$. The next step is to evaluate the posterior conditional probabilities for each Hamiltonian eigenvalue using the standard Born rule, $p(e_{\mathrm{i}}^{j}|\lambda_{\mathrm{i}}^{\alpha})=\text{tr}[\Pi_{\textrm{i}}^{j}\mathrm{P}_{\textrm{i}}^{\alpha}\rho_{\textrm{i}}\mathrm{P}_{\textrm{i}}^{\alpha}/\text{tr}(\mathrm{P}_{\textrm{i}}^{\alpha}\rho_{\textrm{i}})]$ and $p(e_{\mathrm{f}}^{k}|\lambda_{\mathrm{f}}^{\beta})=\text{tr}[\Pi_{\textrm{f}}^{k}\mathrm{P}_{\textrm{f}}^{\beta}\rho_{\textrm{f}}\mathrm{P}_{\textrm{f}}^{\beta}/\text{tr}(P_{\textrm{f}}^{\beta}\rho_{\textrm{f}})]$. Summing over all possible measurement outcomes, one finds lutz_2020 ; par20 ; mic20a ; str20 ; zha22 ; lutz_2023 ; lutz_2024 , $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})=\sum_{\alpha,\beta}p(e_{\mathrm{i}}^{j}|\lambda_{\mathrm{i}}^{\alpha})\,p(e_{\mathrm{f}}^{k}|\lambda_{\mathrm{f}}^{\beta})\,p(\lambda_{\mathrm{i}}^{\alpha},\lambda_{\mathrm{f}}^{\beta})$. The two joint distributions, $p_{\textrm{TPM}}$ and $p_{\textrm{DBN}}$, coincide in the absence of quantum coherence in the energy basis, yielding identical energy statistics. However, because the states $\rho_{\textrm{i},\textrm{f}}$ are diagonal in their eigenbases, measurement backaction is strongly reduced in the DBN protocol in the presence of energy coherences, compared to the TPM approach. Importantly, the Bayesian network distribution $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})$ is accessible experimentally mic20a .
@@
The joint probability $p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})$ in Eq. (1) can be derived for the quantum Otto engine in a similar manner by extending the previous considerations to the four corners of the cycle. The explicit (and lengthy) expressions for the work output distribution $P(w)$ for the two measurement schemes are given in the Supplemental Material. We here indicate the corresponding average work outputs $\langle w\rangle=\int dw\,w\,P(w)$. We have
@@
Remarkably, the average work, $\langle w\rangle_{\mathrm{DBN}}$ in Eq. (3), obtained via Bayesian networks always agrees exactly with the unmeasured result $W$. By contrast, the two-point measurement expression, $\langle w\rangle_{\mathrm{TPM}}$ in Eq. (2), differs from it in general due to measurement backaction. A comparison of Eqs. (2) and (3) reveals that this discrepancy originates from the presence of quantum coherence during the cycle: $\langle w\rangle_{\mathrm{TPM}}=\langle w\rangle_{\mathrm{DBN}}=W$ only if the states of the working substance after the hot and cold isochores are diagonal in the eigenbasis of their respective Hamiltonians, that is, $\tilde{\rho}_{1}=\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})$ after the cold isochore and $\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})=\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})$ after the hot isochore. For the specific Hamiltonian considered here, this is achieved either in the absence of the transverse driving ($g_{i}=0$), or when the system fully thermalizes with the baths ($t_{\mathrm{h},\mathrm{c}}\to\infty$). In these cases the two work distributions (1) also coincide, $P_{\mathrm{TPM}}(w)=P_{\mathrm{DBN}}(w)$. For arbitrary system Hamiltonians $H_{i}$, and finite thermalization times, these conditions can be generically formulated as (Supplemental Material)
$\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\rho)\!=\mathrm{D}_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\rho)\text{ and }\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}(\rho)\!=\mathrm{D}_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}(\rho)\!\!$ (4)
since the output of the generalized amplitude damping map $\mathcal{T}_{t}$ is diagonal only if the input is already diagonal. Equation (4) is satisfied whenever the evolution operators $\mathcal{U}_{\mathrm{k,e}}$ do not induce coherences on incoherent states. It is worth noting that dynamic Bayesian networks are different from so-called collective measurements which have been shown to reduce measurement backaction acin_2017 ; wu19 . Whereas the measured average work in that scheme approaches the unmeasured value, after proper parameter optimization, it does not reach it in general.
@@
Both measurement protocols are unavoidably accompanied by measurement backaction. However, the above findings suggest that the backaction induced by the dynamic Bayesian network scheme is significantly weaker than that of the two-point measurement approach. To quantify the disturbance induced by a measurement sequence, we compare the unmeasured steady state $\tilde{\rho}_{1}$ with the averaged state obtained after a full engine cycle interspersed by five measurements, back to the initial corner of the cycle. We find $\langle\rho\rangle_{\textrm{TPM}}=\mathrm{D}_{\mathrm{e}}\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathrm{D}_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathrm{D}_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})\neq\tilde{\rho}_{1}$ whereas $\langle\rho\rangle_{\textrm{DBN}}=\tilde{\rho}_{1}$. The Bayesian network approach therefore does not perturb the state of the working substance on average. In this sense, it is minimally invasive, in sharp contrast to the two-point measurement protocol. This difference arises because, in the Bayesian network scheme, measurements are performed in the eigenbasis of the density operator, in which the state is trivially diagonal. Figure 2 displays the trace distance, $d(\langle\rho\rangle,\tilde{\rho}_{1})$ nie10 , between the averaged measured state and the unmeasured state, as a function of the thermalization rate $\lambda$. These results indicate that two-point measurement experiments become increasingly invasive as the amount of coherence generated during the engine cycle grows for smaller $\lambda$.
@@
Equation (5) was derived assuming full thermalization and in the absence of quantum coherence agarwalla_2021_2 ; agarwalla_2021 ; das23 ; moh23 ; moh23a ; watanabe_2025 . Figure 4 reveals that the bound is actually violated in the presence of quantum coherence, and this for both measurement schemes, for finite relaxation times. Coherence is seen to augment work output fluctuations. We stress that the value of the transverse driving strength is here one order of magnitude smaller than the frequencies to ensure that the engine runs in the quasistatic regime.
@@ I.2 Dynamic Bayesian network approach (DBN)
In the case of the DBN protocol lutz_2020s ; par20s ; mic20as ; str20s ; zha22s ; lutz_2023s ; lutz_2024s , it is convenient to consider five copies of the working substance lutz_2023s : $\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}$ (If only $p_{w}$ and $p_{q_{\mathrm{h}}}$ are required, only four copies are necessary). Given that the quantum engine reaches a steady state after a few cycles, in order to obtain the five copies, it is enough to initialize the engine with five arbitrary states $\rho$ and wait until they all reach the steady state $\tilde{\rho}_{1}$. Let $\tilde{\rho}_{i}=\sum_{\alpha}\lambda_{i}^{\alpha}\mathrm{P}_{i}^{\alpha}$ be the spectral decomposition of the state of the working substance at corner $i$ of the engine cycle. The measurement protocol to obtain $p^{\mathrm{DBN}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ proceeds as follows:
@@
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\otimes\tilde{\rho}_{2}\otimes\tilde{\rho}_{2}\otimes\tilde{\rho}_{2}.$ (S18)
@@
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathrm{P}_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\mathrm{P}_{2}^{\beta}\otimes\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}/\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})]\otimes\tilde{\rho}_{3}\otimes\tilde{\rho}_{3}.$ (S19)
@@
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathrm{P}_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\mathrm{P}_{2}^{\beta}\otimes\mathrm{P}_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}/\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})]\mathrm{P}_{3}^{\gamma}\otimes\mathcal{U}_{\mathrm{e}}[\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}/\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})]\otimes\tilde{\rho}_{4}.$ (S20)
@@
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathrm{P}_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\mathrm{P}_{2}^{\beta}\otimes\mathrm{P}_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}/\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})]\mathrm{P}_{3}^{\gamma}$ (S21)
    $\otimes\mathrm{P}_{4}^{\delta}\mathcal{U}_{\mathrm{e}}[\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}/\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})]\mathrm{P}_{4}^{\delta}\otimes\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}[\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}/\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})].$
@@
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathrm{P}_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\mathrm{P}_{2}^{\beta}\otimes\mathrm{P}_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}/\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})]\mathrm{P}_{3}^{\gamma}$ (S22)
    $\otimes\mathrm{P}_{4}^{\delta}\mathcal{U}_{\mathrm{e}}[\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}/\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})]\mathrm{P}_{4}^{\delta}\otimes\mathrm{P}_{1}^{\epsilon}\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}[\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}/\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})]\mathrm{P}_{1}^{\epsilon}.$
@@
$p(\lambda_{1}^{\alpha},\lambda_{2}^{\beta},\lambda_{3}^{\gamma},\lambda_{4}^{\delta},\lambda_{1}^{\epsilon})$ $=\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}$ (S23)
$\times\text{tr}\left\{\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left[\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]\right\}.$